# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ E.1 Derivation of $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Eq. (100) provides an expression for $p_{f_{0}}\left({x}\right)$. However, computing the integral on the right-hand side (r.h.s.) is not straightforward. To circumvent this, we introduce a second approach that is asymptotically exact in the limit of $N_{1}\rightarrow\infty$.
In App. D.3, we derived expressions for $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ for different cases. These distributions are generally Gaussian, with cumulants dependent on $S_{\boldsymbol{w}_{c}}\equiv\sum_{j}^{N_{1}}w^{(1)}_{cj}$ and $S_{\boldsymbol{w}_{c}^{2}}\equiv\sum_{j}^{N_{1}}\left(w^{(1)}_{cj}\right)^{2}$ . The standard deviations can be viewed as estimates for the magnitude of fluctuations from the mean value.
@@
$p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\mu\left(S_{\boldsymbol{w}_{c}}\right),\sigma^{2}\left(S_{\boldsymbol{w}_{c}^{2}}\right)\right),$ (103)
@@
Since the distribution of $\mu\left(S_{\boldsymbol{w}_{c}}\right)$ does not narrow asymptotically around a deterministic value, $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ will differ for various nodes, $\{O^{(c)}\}$ , even in the limit of infinite size. We refer to the emergence of this discrepancy in distributions among nodes of the same layer as Node Symmetry Breaking (NSB). Section 4 elucidates how NSB and IGB are interrelated phenomena, with IGB being a direct consequence of this symmetry breaking. In Sec. 5.2, we will delve deeper into NSB, illustrating how this phenomenon manifests not only in output layers but also in intermediate hidden layers of deep architectures.
Returning to the initial problem, let’s consider a given initialization fixing the configuration $\mathcal{W}$, and in particular, the two random vectors $w^{(1)}_{0\;\boldsymbol{\cdot}}$ and $w^{(1)}_{1\;\boldsymbol{\cdot}}$ . This leads to two corresponding distributions for $p_{O^{(0)}}^{(\chi)}\left({x}\right)$ and $p_{O^{(1)}}^{(\chi)}\left({x}\right)$ . To calculate the probability that $O^{(0)}>O^{(1)}$ , we first transition to the distribution $p_{\Delta_{O}}^{(\chi)}\left({x}\right)$ . Since $\Delta_{O}$ is the difference between two Gaussian random variables, it follows that:
$p_{\Delta_{O}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\mu\left(S_{\boldsymbol{w}_{0}}\right)-\mu\left(S_{\boldsymbol{w}_{1}}\right),\sigma^{2}\left(S_{\boldsymbol{w}_{0}^{2}}\right)+\sigma^{2}\left(S_{\boldsymbol{w}_{1}^{2}}\right)\right).$ (104)
@@
$p_{\Delta_{O}}^{(\chi)}\left({x}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(x;\mu\left(S_{\boldsymbol{w}_{0}}\right)-\mu\left(S_{\boldsymbol{w}_{1}}\right),2\sigma^{2}_{\infty}\right).$ (105)
@@
$f_{0}\left(\mathcal{W}\right)=\int_{0}^{\infty}\mathcal{N}\left(y;\Delta_{\mu}(f_{0}),2\sigma^{2}_{\infty}\right)dy.$ (107)
Given the centers of the two Gaussian distributions $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , specifically their difference $\Delta_{\mu}$, Eq. (107) provides the corresponding value of $f_{0}$. Inverting Eq. (107) numerically yields $\Delta_{\mu}(f_{0})$ associated with a given value of $f_{0}$.
Notably, $\mu\left(S_{\boldsymbol{w}_{c}}\right)\sim p_{\mu_{c}}\left({x}\right)$ is a r.v.; consequently, $\Delta_{\mu}(f_{0})$ is a r.v. as well. From the monotonicity of the relationship $\Delta_{\mu}(f_{0})$, it follows that
$p_{f_{0}}\left({x}\right)\,dx=p_{\Delta_{\mu}(f_{0})}\left({\Delta_{\mu}(x)}\right)\,d\Delta_{\mu}$ (108)
@@
Given that $p_{\mu_{c}}\left({x}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(x;0,\hat{\sigma}^{2}_{\infty}\right)$ , [^13] The quantity $\hat{\sigma}^{2}_{\infty}$ is model-specific; for more details, see App. D.2. Eq. (109) represents an integral of a product of two independent Gaussian distributions. [^14] The product of Gaussian distributions itself is proportional to a Gaussian distribution. Solving this integral gives us
$p_{\Delta_{\mu}(f_{0})}\left({\Delta_{\mu}(x)}\right)=\mathcal{N}\left(\Delta_{\mu}(x);0,2\hat{\sigma}^{2}_{\infty}\right).$ (110)
Remark 2.
While Eq. (110) illustrates a Gaussian distribution in terms of the variable $\Delta_{\mu}(f_{0})$, it is important to note that $p_{f_{0}}\left({x}\right)$ will not generally be Gaussian since $\Delta_{\mu}\left(f_{0}\right)$ is a non-linear function.
max pool peaks: A practical example for computing $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Utilizing the results from App. E.1, we can calculate the theoretical prediction for $p_{f_{0}}\left({x}\right)$, particularly focusing on the probability density at extreme peaks empirically observed with max pooling.
Suppose we have a dataset of $D$ elements, and we aim to calculate the probability density $p_{f_{0}}\left({0}\right)$. The steps involved are:
  • Computing $\Delta^{(T)}$: First, we need to invert Eq. (107):
    $f_{0}=1-\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{\Delta_{\mu}(f_{0})}{2\sigma_{\infty}}\right)\right)\Rightarrow\Delta_{\mu}(f_{0})=2\sigma_{\infty}\operatorname{erf}^{-1}\left(2f_{0}-1\right).$ (111)
    $\Delta^{(T)}$ is the threshold value; for $\Delta_{\mu}(f_{0})<\Delta^{(T)}$, Eq. (107) yields a value below $\frac{1}{D}$, corresponding to $f_{0}=0$ due to the dataset size $D$ resolution.
  • Integrating Eq. (110): Next, we integrate over all $\Delta_{\mu}(f_{0})<\Delta^{(T)}$:
@@
Generalizing for arbitrary probability mass: For the probability mass between $f_{0}^{(min)}$ and $f_{0}^{(max)}$, the computation is:
[…]
These computations facilitate a precise understanding of the impact of max pooling on $f_{0}$, especially in terms of evaluating probability densities at the extreme peaks.
E.2 Beyond $\boldsymbol{f_{0}}$: relating the confidence in model assignments to IGB
While our discussion is based on an intuitive and clearly interpretable measure, namely $f_{0}$ and its distribution, this measure has a significant limitation. $f_{0}$ indicates the fraction of the dataset assigned to class 0 by the network at initialization, but it doesn’t provide information about the average confidence with which this assignment is made for a given dataset element.
It is important to note that, although $f_{0}$ is a central focus of our narrative for its intuitive clarity, it is not the main focus of our analysis. Our analyses are instead centered around the output nodes, from which we derive distributions, and only subsequently, through these and Eq. (3), we transition to the variable $f_{0}$.
@@ E.2 Beyond $\boldsymbol{f_{0}}$: relating the confidence in model assignments to IGB
$O^{(c)}\rightarrow C_{O^{(c)}}\equiv\frac{e^{O^{(c)}}}{\sum_{i}e^{O^{(i)}}}\,.$ (114)
@@
$R_{C}\equiv\frac{C_{O^{(0)}}}{C_{O^{(1)}}}=e^{\Delta_{O}}\,$, (116)
where we remind the reader that $\Delta_{O}\equiv O^{(0)}-O^{(1)}$.
Since we have shown that the distribution of a generic output, $p_{O^{(c)}}^{(\chi)}\left({x}\right)$, is Gaussian (see Eq. (58)), it follows that
$p_{\Delta_{O}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\Delta_{\mu},2\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right)\,$, (117)
@@
$p_{R_{C}}^{(\chi)}\left({x}\right)=\frac{1}{\Delta_{\mu}\sqrt{4\pi\textrm{Var}_{\chi}\left({O^{(c)}}\right)}}e^{-\frac{(\ln(x)-\Delta_{\mu})^{2}}{4\textrm{Var}_{\chi}\left({O^{(c)}}\right)}}.$ (118)
The first two cumulants (over the distribution of the data) of $p_{R_{C}}^{(\chi)}\left({x}\right)$ are:
$\left\langle{R_{C}}\right\rangle$ $=e^{\Delta_{\mu}+\textrm{Var}_{\chi}\left({O^{(c)}}\right)}\,$, (119)
$\textrm{Var}_{\chi}\left({R_{C}}\right)$ $=e^{2\Delta_{\mu}}\left(e^{2\textrm{Var}_{\chi}\left({O^{(c)}}\right)}-1\right)e^{2\textrm{Var}_{\chi}\left({O^{(c)}}\right)}\,.$ (120)
@@
$\lim_{\Delta_{\mu}\to 0}\penalty 10000\ \lim_{\textrm{Var}_{\chi}\left({O^{(c)}}\right)\to 0}\left\langle{R_{C}}\right\rangle$ $=1\,$, (121)
$\lim_{\textrm{Var}_{\chi}\left({O^{(c)}}\right)\to 0}\textrm{Var}_{\chi}\left({R_{C}}\right)$ $=0\,.$ (122)
@@
We may encounter a scenario where $\Delta_{\mu}=0$ and $\textrm{Var}_{\chi}\left({O^{(c)}}\right)\neq 0$. Even though this indicates an absence of IGB, the distribution for $R_{C}$ is non-degenerate. Like the previous case, the dataset will be evenly divided in assignments between the different classes. However, the typical confidence is different for each class (it is described by Eq. (118)).
@@
In conclusion, the approach shown in this section paves the way for more accurate analyses. For example, knowing the $\{C_{O^{(c)}}\}$, we could calculate the entropy $H\left(\{C_{O^{(c)}}\}\mid\mathcal{W},\chi\right)$ and reformulate the absence of IGB as a principle of maximum entropy. However, it is important to note that the presence of the normalization factor complicates the derivation of the distribution of $C_{O^{(c)}}$, which can be avoided by using $R_{C}$. This also provides a natural variable to encapsulate the gap between the assignment variables. Finally, although we focused on the binary problem for simplicity, extensions to the multi-class case are possible using approaches similar to those described in App. H, for example by considering the ratio of the two classes with the highest confidence.
@@ Appendix F Conditions for the emergence/absence of IGB
In our exploration of how activation functions can induce IGB, we have focused on specific, emblematic examples, such as the linear activation function and the ReLU. These were chosen for their analytical tractability. However, the underlying principles extend beyond these cases. In this section, we’ll broaden the discussion, classifying activation functions into distinct categories based on their influence on IGB.
@@
This symmetry breaking is intricately linked to how the ReLU, applied at the first hidden layer, transforms null-averaged inputs into random variables, $g^{(1)}_{i}$, with a non-zero mean.
To generalize our findings, we’ll now examine a typical activation function from each category and demonstrate how the presence or absence of IGB manifests. Our analysis will maintain the same foundational assumptions as before: Gaussian-distributed i.i.d. inputs and Kaiming initialization for the weights. As established in App. D.2, regardless of the activation function, we observe that:
$p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\right).$ (123)

### v5

@@ E.1 Derivation of $\boldsymbol{f_{G_{0}}\left({g_{0}}\right)}$
Eq. (100) provides an expression for $f_{G_{0}}\left({g_{0}}\right)$. However, computing the integral on the right-hand side (r.h.s.) is not straightforward. To circumvent this, we introduce a second approach that is asymptotically exact in the limit of $N_{1}\rightarrow\infty$.
In App. D.3, we derived expressions for $f_{O_{c}}^{(\chi)}\left({o}\right)$ for different cases. These distributions are generally Gaussian, with cumulants dependent on $S_{\boldsymbol{w}_{c}}\equiv\sum_{j}^{N_{1}}w^{(1)}_{cj}$ and $S_{\boldsymbol{w}_{c}^{2}}\equiv\sum_{j}^{N_{1}}\left(w^{(1)}_{cj}\right)^{2}$ . The standard deviations can be viewed as estimates for the magnitude of fluctuations from the mean value.
@@
$f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;\mu\left(S_{\boldsymbol{w}_{c}}\right),\sigma^{2}\left(S_{\boldsymbol{w}_{c}^{2}}\right)\right),$ (103)
@@
Since the distribution of $\mu\left(S_{\boldsymbol{w}_{c}}\right)$ does not narrow asymptotically around a deterministic value, $f_{O_{c}}^{(\chi)}\left({o}\right)$ will differ for various nodes, $\{O_{c}\}$ , even in the limit of infinite size. We refer to the emergence of this discrepancy in distributions among nodes of the same layer as Node Symmetry Breaking (NSB). Section 4 elucidates how NSB and IGB are interrelated phenomena, with IGB being a direct consequence of this symmetry breaking. In Sec. 5.2, we will delve deeper into NSB, illustrating how this phenomenon manifests not only in output layers but also in intermediate hidden layers of deep architectures.
Returning to the initial problem, let’s consider a given initialization fixing the configuration $\mathcal{W}$, and in particular, the two random vectors $w^{(1)}_{0\;\boldsymbol{\cdot}}$ and $w^{(1)}_{1\;\boldsymbol{\cdot}}$ . This leads to two corresponding distributions for $f_{O_{0}}^{(\chi)}\left({o}\right)$ and $f_{O_{1}}^{(\chi)}\left({o}\right)$ . To calculate the probability that $O_{0}>O_{1}$ , we first transition to the distribution $f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)$ . Since $\Delta_{O}$ is the difference between two Gaussian random variables, it follows that:
$f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)=\mathcal{N}\left(\delta_{o};\mu\left(S_{\boldsymbol{w}_{0}}\right)-\mu\left(S_{\boldsymbol{w}_{1}}\right),\sigma^{2}\left(S_{\boldsymbol{w}_{0}^{2}}\right)+\sigma^{2}\left(S_{\boldsymbol{w}_{1}^{2}}\right)\right).$ (104)
@@
$f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(\delta_{o};\mu\left(S_{\boldsymbol{w}_{0}}\right)-\mu\left(S_{\boldsymbol{w}_{1}}\right),2\sigma^{2}_{\infty}\right).$ (105)
@@
$G_{0}\left(\mathcal{W}\right)=\int_{0}^{\infty}\mathcal{N}\left(\delta_{o};\Delta_{\mu}(G_{0}),2\sigma^{2}_{\infty}\right)d\delta_{o}.$ (107)
Given the centers of the two Gaussian distributions $f_{O_{c}}^{(\chi)}\left({o}\right)$ , specifically their difference $\Delta_{\mu}$, Eq. (107) provides the corresponding value of $G_{0}$. Inverting Eq. (107) numerically yields $\Delta_{\mu}(G_{0})$ associated with a given value of $G_{0}$.
Notably, $\mu\left(S_{\boldsymbol{w}_{c}}\right)\sim f_{\mu_{c}}\left({m}\right)$ is a r.v.; consequently, $\Delta_{\mu}(G_{0})$ is a r.v. as well. From the monotonicity of the relationship $\Delta_{\mu}(G_{0})$, it follows that
$f_{G_{0}}\left({g_{0}}\right)\,dg_{0}=f_{\Delta_{\mu}(G_{0})}\left({\delta_{m}(g_{0})}\right)\,d\delta_{m}$ (108)
@@
$\begin{split}f_{\Delta_{\mu}(G_{0})}\left({\delta_{m}(g_{0})}\right)=\int_{-\infty}^{\infty}f_{\mu_{1}}\left({m}\right)f_{\mu_{0}}\left({m+\delta_{m}(g_{0})}\right)dm=\int_{-\infty}^{\infty}f_{\mu_{c}}\left({m}\right)f_{\mu_{c}}\left({m+\delta_{m}(g_{0})}\right)dm,\end{split}$ (109)
Given that $f_{\mu_{c}}\left({m}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(m;0,\hat{\sigma}^{2}_{\infty}\right)$ , [^13] The quantity $\hat{\sigma}^{2}_{\infty}$ is model-specific; for more details, see App. D.2. Eq. (109) represents an integral of a product of two independent Gaussian distributions. [^14] The product of Gaussian distributions itself is proportional to a Gaussian distribution. Solving this integral gives us
$f_{\Delta_{\mu}(G_{0})}\left({\delta_{m}(g_{0})}\right)=\mathcal{N}\left(\delta_{m}(g_{0});0,2\hat{\sigma}^{2}_{\infty}\right).$ (110)
Remark 2.
While Eq. (110) illustrates a Gaussian distribution in terms of the variable $\Delta_{\mu}(G_{0})$, it is important to note that $f_{G_{0}}\left({g_{0}}\right)$ will not generally be Gaussian since $\Delta_{\mu}\left(G_{0}\right)$ is a non-linear function.
max pool peaks: A practical example for computing $\boldsymbol{f_{G_{0}}\left({g_{0}}\right)}$
Utilizing the results from App. E.1, we can calculate the theoretical prediction for $f_{G_{0}}\left({g_{0}}\right)$, particularly focusing on the probability density at extreme peaks empirically observed with max pooling.
Suppose we have a dataset of $D$ elements, and we aim to calculate the probability density $f_{G_{0}}\left({0}\right)$. The steps involved are:
  • Computing $\Delta^{(T)}$: First, we need to invert Eq. (107):
    $g_{0}=1-\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{\Delta_{\mu}(g_{0})}{2\sigma_{\infty}}\right)\right)\Rightarrow\Delta_{\mu}(g_{0})=2\sigma_{\infty}\operatorname{erf}^{-1}\left(2g_{0}-1\right).$ (111)
    $\Delta^{(T)}$ is the threshold value; for $\Delta_{\mu}(g_{0})<\Delta^{(T)}$, Eq. (107) yields a value below $\frac{1}{D}$, corresponding to $g_{0}=0$ due to the dataset size $D$ resolution.
  • Integrating Eq. (110): Next, we integrate over all $\Delta_{\mu}(g_{0})<\Delta^{(T)}$:
@@
Generalizing for arbitrary probability mass: For the probability mass between $g_{0}^{(min)}$ and $g_{0}^{(max)}$, the computation is:
@@
These computations facilitate a precise understanding of the impact of max pooling on $G_{0}$, especially in terms of evaluating probability densities at the extreme peaks.
E.2 Beyond $\boldsymbol{G_{0}}$: relating the confidence in model assignments to IGB
While our discussion is based on an intuitive and clearly interpretable measure, namely $G_{0}$ and its distribution, this measure has a significant limitation. $G_{0}$ indicates the fraction of the dataset assigned to class 0 by the network at initialization, but it doesn’t provide information about the average confidence with which this assignment is made for a given dataset element.
It is important to note that, although $G_{0}$ is a central focus of our narrative for its intuitive clarity, it is not the main focus of our analysis. Our analyses are instead centered around the output nodes, from which we derive distributions, and only subsequently, through these and Eq. (3), we transition to the variable $G_{0}$.
@@ E.2 Beyond $\boldsymbol{G_{0}}$: relating the confidence in model assignments to IGB
$O_{c}\rightarrow C_{o_{c}}\equiv\frac{e^{O_{c}}}{\sum_{i}e^{o_{i}}}\,.$ (114)
@@
$R_{C}\equiv\frac{C_{o_{0}}}{C_{o_{1}}}=e^{\Delta_{O}}\,$, (116)
where we remind the reader that $\Delta_{O}\equiv O_{0}-O_{1}$.
Since we have shown that the distribution of a generic output, $f_{O_{c}}^{(\chi)}\left({o}\right)$, is Gaussian (see Eq. (58)), it follows that
$f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)=\mathcal{N}\left(\delta_{o};\Delta_{\mu},2\textrm{Var}_{\chi}\left({O_{c}}\right)\right)\,$, (117)
@@
$f_{R_{C}}^{(\chi)}\left({r}\right)=\frac{1}{\Delta_{\mu}\sqrt{4\pi\textrm{Var}_{\chi}\left({O_{c}}\right)}}e^{-\frac{(\ln(r)-\Delta_{\mu})^{2}}{4\textrm{Var}_{\chi}\left({O_{c}}\right)}}.$ (118)
The first two cumulants (over the distribution of the data) of $f_{R_{C}}^{(\chi)}\left({r}\right)$ are:
$\left\langle{R_{C}}\right\rangle$ $=e^{\Delta_{\mu}+\textrm{Var}_{\chi}\left({O_{c}}\right)}\,$, (119)
$\textrm{Var}_{\chi}\left({R_{C}}\right)$ $=e^{2\Delta_{\mu}}\left(e^{2\textrm{Var}_{\chi}\left({O_{c}}\right)}-1\right)e^{2\textrm{Var}_{\chi}\left({O_{c}}\right)}\,.$ (120)
@@
$\lim_{\Delta_{\mu}\to 0}\penalty 10000\ \lim_{\textrm{Var}_{\chi}\left({O_{c}}\right)\to 0}\left\langle{R_{C}}\right\rangle$ $=1\,$, (121)
$\lim_{\textrm{Var}_{\chi}\left({O_{c}}\right)\to 0}\textrm{Var}_{\chi}\left({R_{C}}\right)$ $=0\,.$ (122)
@@
We may encounter a scenario where $\Delta_{\mu}=0$ and $\textrm{Var}_{\chi}\left({O_{c}}\right)\neq 0$. Even though this indicates an absence of IGB, the distribution for $R_{C}$ is non-degenerate. Like the previous case, the dataset will be evenly divided in assignments between the different classes. However, the typical confidence is different for each class (it is described by Eq. (118)).
@@
In conclusion, the approach shown in this section paves the way for more accurate analyses. For example, knowing the $\{C_{o_{c}}\}$, we could calculate the entropy $H\left(\{C_{o_{c}}\}\mid\mathcal{W},\chi\right)$ and reformulate the absence of IGB as a principle of maximum entropy. However, it is important to note that the presence of the normalization factor complicates the derivation of the distribution of $C_{o_{c}}$, which can be avoided by using $R_{C}$. This also provides a natural variable to encapsulate the gap between the assignment variables. Finally, although we focused on the binary problem for simplicity, extensions to the multi-class case are possible using approaches similar to those described in App. H, for example by considering the ratio of the two classes with the highest confidence.
@@ Appendix F Conditions for the emergence/absence of IGB
In our exploration of how activation functions can induce IGB, we have focused on specific, emblematic examples, such as the linear activation function and the ReLU. These were chosen for their analytical tractability. However, the underlying principles extend beyond these cases. In this section, we’$l$ broaden the discussion, classifying activation functions into distinct categories based on their influence on IGB.
@@
This symmetry breaking is intricately linked to how the ReLU, applied at the first hidden layer, transforms null-averaged inputs into random variables, $H^{(1)}_{i}$, with a non-zero mean.
To generalize our findings, we’$l$ now examine a typical activation function from each category and demonstrate how the presence or absence of IGB manifests. Our analysis will maintain the same foundational assumptions as before: Gaussian-distributed i.i.d. inputs and Kaiming initialization for the weights. As established in App. D.2, regardless of the activation function, we observe that:
$f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}\left(z{}{};0,\sigma^{2}_{w}\right).$ (123)